40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 12 22:51:26 2016, Koji, Update, General, Clipping could not be improved by the PZTs - NEED ANOTHER VENT 
    Reply  Thu Oct 13 00:23:54 2016, ericq, Update, VAC, Pre-vent prep 
       Reply  Thu Oct 13 09:30:38 2016, Steve, Update, General, vent 79 is happening vent79precond.pngventing.png
          Reply  Thu Oct 13 16:48:34 2016, Steve, Update, VAC, vent 79 is completed vent_79.png
             Reply  Thu Oct 13 19:28:14 2016, gautam, Update, General, in-air alignment 
                Reply  Fri Oct 14 10:31:53 2016, gautam, Update, General, doors are off ITMY and BS/PRM chambers 
                   Reply  Fri Oct 14 18:33:55 2016, gautam, Update, General, AS clipping investigations 7x
                      Reply  Mon Oct 17 22:45:16 2016, gautam, Update, General, AS beam centered on all OMs 7x
                         Reply  Tue Oct 18 18:56:57 2016, gautam, Update, General, OM5 rotated to bypass OMC, green scatter is from window to PSL table OMCchamber.pdfgreenGhosts.JPGIMG_3322.JPGIMG_3326.JPG
                            Reply  Wed Oct 19 16:41:55 2016, gautam, Update, General, Heavy doors back on 8x
             Reply  Thu Oct 20 09:17:08 2016, Steve, Update, VAC, Pumpdown 80 has started readyPd80.pngbeforePd80.pngpd80at1h.png7d_atm.png
                Reply  Thu Oct 20 15:59:52 2016, Steve, Update, VAC, Pumpdown 80 has completed 7hPd80.png8hPd80.png9hPd80.png
                   Reply  Fri Oct 21 02:06:20 2016, gautam, Update, General, IFO recovery 
                      Reply  Mon Oct 24 11:39:13 2016, gautam, Update, General, ALS recovered ALSOutOfLoop20161024.pdfXendPDHOLTF20161024.pdf
                   Reply  Fri Oct 21 09:28:21 2016, Steve, Update, VAC, Vac Normal reached VacNormal.png
                      Reply  Wed Nov 23 11:32:13 2016, Steve, Update, VAC, TP3 drypump replaced again 
                         Reply  Fri Nov 10 08:52:21 2017, Steve, Update, VAC, TP3 drypump replaced 
                            Reply  Fri Feb 16 15:57:58 2018, Steve, Update, VAC, TP3 drypump replaced 
                Reply  Thu Jul 12 16:18:22 2018, Steve, Update, VAC, Vent preparations before_Vent.pngbefore_Vent_cond.png
                   Reply  Thu Jul 12 23:59:14 2018, gautam, Update, General, Vent objectives and prep 
                      Reply  Fri Jul 13 02:52:11 2018, gautam, Update, General, Vent objectives and prep preVentStatus.png
                         Reply  Fri Jul 13 17:13:45 2018, Koji, Update, General, Vent objectives and prep P_20180713_020603.jpg
                         Reply  Fri Jul 13 18:01:13 2018, gautam, Update, General, Low power MC 
Message ID: 12555     Entry time: Wed Oct 12 22:51:26 2016     Reply to this: 12556
Author: Koji 
Type: Update 
Category: General 
Subject: Clipping could not be improved by the PZTs - NEED ANOTHER VENT 

[Gautam Koji]

We engaged the HV driver to the output port PZTs, hoping to mitigate the AS port clipping. Basically, the range of the PZT is not enough to make the beam look clean. Also, our observation suggested there are possible multiple clipping in the chamber. We need another vent to make the things clearly right. Eric came in the lab and preparing the IFO for it.


1. Before the test, the test masses have been aligned with the dither servo.

2. We looked at the beam shape on the AS camera with a single bounce beam. We confirmed that the beam is hard-clipped at the upper and left sides of the beam on the video display. This clipping is not happening outside of the chamber.

3. We brought an HV power supply to the short OMC rack. There is a power supply cable with two spades. The red and black wires are +150V and GND respectively.

4. The voltage of +/-10V was applied on each of the four PZT drive inputs. We found that the motion of the beam on the camera is tiny and in any case, we could not improve the beam shape.

5. We wondered that if we are observing ANY improvement of the clipping. For this purpose, we aligned AS110 sensor every time we gave the misalignment with the PZTs. Basically, we are at the alignment to have the best power we can get. We thought this was weird.

6. Then we moved the AS port spot with the ITMX. We could clearly make the spot more round. However, this reduced the power at the AS port reduced by ~15%. When the beam was further clipped, the power went down again. Basically, the initial alignment gave us the max power we could get. As the max power was given with the clipped beam, we get confused and feel safer to check the situation with the chambers open.

During this investigation, we moved the AS port opitcs and the AS camera. So they are not too precise reference of the alignment. The PZT HV setup has been removed.

ELOG V3.1.3-